OpenStudy (anonymous):

Subtract: x-3 x-4 --- - --- 3 4

OpenStudy (anonymous):

[4(x-3) - 3(x-4)]/12 (4x-12-3x-12)/12 x-24/12

OpenStudy (anonymous):

\[\frac{x-3}{3}-\frac{x-4}{4}=\frac{4(x-3)-3(x-4)}{12}\]

OpenStudy (anonymous):

clean up numerator get matematicalpro answer